radomes.openskiesproject.org
Davante69 Davante69
  • 19-10-2014
  • Mathematics
Answered

All that is left in a packet of candy are 7 reds, 4 greens, and 2 blues, What is the probability that a random drawing yields a green followed by a red assuming that the first candy drawn is put back into the packet? 

Verified Answer :

jer1449 jer1449
  • 20-10-2014

probability of drawing a green = 4/13
Probability of drawing a red = 2/13
probability of drawing a green and then a red is 4/13 * 2/13 = 6/169

Answer Link
scorp111786 scorp111786
  • 20-10-2014

Multiply 7(reds) by 4 (greens) to get 2/13

Answer Link

Related Questions

which function has real zeros at x 3 and x 7 fx x2 4x 21 fx x2 4x 21 fx x2 10x 21 fx x2 10x 21
The main purpose of a balance sheet is to comply with IRS laws obtain a bank loan show a businesss financial state show profits earned
what is one of the three things the opponents of the political machines argued about
what were the two names of the rivers that were in southwest asia
I need help with 14
The two main types of fermentation are called
When is it appropriate to break the rule that requires subjects joined by and to use a plural verb
write 3010 as a whole number
What is the other meaning for inanimates
Which equation represents a line that passes through 2 12 and has a slope of 3